OpenStudy (anonymous):

MEDAL AND FAN The variable t represents the temperature at 6 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Washington dropped 19°F between 6 p.m. and 10 p.m. The temperature at 10 p.m. was –12°F. What was the temperature at 6 p.m.?

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A. 12 – 19 = t B. 19 + 12 = t C. t – 19 = –12 D. 19 – t = –12

OpenStudy (anonymous):

Idk if it shows it or not...

OpenStudy (anonymous):

It doesn't. Can you type it, please?

OpenStudy (anonymous):

Wait one sec, sorry!

OpenStudy (anonymous):

Don't worry, its fine!

OpenStudy (anonymous):

Here! It'll probably be better this way. :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

C would be the answer

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You're welcome! Thank you for the medal!